OpenStudy (anonymous):

y=6-x and y=x-2 (2,4) (-4,2) (4,2) no solution

OpenStudy (anonymous):

@Bookworm14

ganeshie8 (ganeshie8):

set the right hand sides equal to each other and solve x

ganeshie8 (ganeshie8):

y=6-x and y=x-2 6-x = x - 2 x = ?

OpenStudy (anonymous):

(4,2)

OpenStudy (bookworm14):

substitute your numbers for xand y , all ordered pairs are wrote in (x,y) format

OpenStudy (anonymous):

am i right

ganeshie8 (ganeshie8):

you're right, but how u got (4, 2) ?

OpenStudy (anonymous):

lining up the lines thanks

ganeshie8 (ganeshie8):

lol whats that